Problem 5-22 CVP Applications; C。ntribution Margin Ratio; Break-Even Analysis; Cost Structure L05-1, し05.3, LO5-4, LO5-5, LOS-6] Due to erratic sales of its sole product-a high-capacity battery for laptop computers-PEM, Inc., has been experiencing financial difficulty for some time. The companys contribution format income statement for the most recent month is given below: Sales (13,500 units x $20 per unit) Variable expenses Contribution margin Fixed expenses Net operating loss $ 270,000 135,000 135,000 150,000 $ (15,000) Required: 1. Compute the companys CM ratio and its break-even point in unit sales and dollar sales 2. The president believes that a $6,300 increase in the monthly advertising budget, combined with an intensified effort by the sales staff, will result in an $86,000 increase in monthly sales. If the president is right, what will be the increase (decrease) in the companys monthly net operating income? 3. Refer to the original data. The sales manager is convinced that a 10% reduction in the selling price, combined with an increase of $31,000 in the monthly advertising budget, will double unit sales. If the sales manager is right, what will be the revised net operating income (loss)? 4. Refer to the original data. The Marketing Department thinks that a fancy new package for the laptop computer battery would grow sales. The new package would increase packaging costs by 0.40 cents per unit. Assuming no other changes, how many units would have to be sold each month to attain a target profit of $4,300? 5. Refer to the original data. By automating, the company could reduce variable expenses by $3 per unit. However, fixed expenses would increase by $56,000 each month a. Compute the new CM ratio and the new break-even point in unit sales and dollar sales b. Assume that the company expects to sell 20,300 units next month. Prepare two contribution format income statements, one assuming that operations are not automated and one assuming that they are. (Show data on a per unit and percentage basis, as well as in total, for each alternative.) C. Would you recommend that the company automate its operations (Assuming that the company expects to sell 20,300)? Complete this question by entering your answers in the tabs below. Req 1 Req 2 Req 3 Req 4 Req 5A Req 5B Req 5C Compute the companys CM ratio and its break-even point in unit sales and dollar sales. (Do not round intermediate calculations.) CM ratio Break-even point in unit sales Break-even point in dollar salesReq 1 Req 2 Req 3 Req 4 Req 5A Req 5B Req 5C The president believes that a $6,300 increase in the monthly advertising budget, combined with an intensified effort by the sales staff, will result in an $86,000 increase in monthly sales. If the president is right, what will be the increase (decrease) in the companys monthly net operating income? (Do not round intermediate calculations.) byReq 1 Req 2 Req 3 Req 4 Req 5A Req 5B Req 5C Refer to the original data. The Marketing Department thinks that a fancy new package for the laptop computer battery would grow sales. The new package would increase packaging costs by 40 cents per unit. Assuming no other changes, how many units would have to be sold each month to attain a target profit of $4,300? (Do not round intermediate calculations. Round final answer to the nearest whole unit.) Show less Unit sales to attain target profitReq 1 Req 2 Req 3 Req 4 Req 5A Req 5B Req 5C Refer to the original data. By automating, the company could reduce variable expenses by $3 per unit. However, fixed expenses would increase by $56,000 each month. Compute the new CM ratio and the new break-even point in unit sales and dollar sales. (Round CM ratio to the nearest whole percent and other answers to the nearest whole number.) CM ratio Break-even point in unit sales Break-even point in dollar salesReq 1 Req 2 Req 3 Req 4 Req 5A Req 5B Req 5C Refer to the original data. By automating, the company could reduce variable expenses by $3 per unit. However, fixed expenses would increase by $56,000 each month. Assume that the company expects to sell 20,300 units next month. Prepare two contribution format income statements, one assuming that operations are not automated and one assuming that they are. (Show data on a per unit and percentage basis, as well as in total, for each alternative.) (Do not round your intermediate calculations. Round your percentage answers to the nearest whole number.) Show less PEM, Inc. Contribution Income Statement Not Automated Automated Total Per Unit Total Per UnitReq 1 Req 2 Req 3 Req 4 Req 5A Req 5B Req 5C Refer to the original data. By automating, the company could reduce variable expenses by $3 per unit. However, fixed expenses would increase by $56,000 each month. Would you recommend that the company automate its operations (Assuming that the company expects to sell 20,300)? Yes ONo
Answer #1

1.CM ratio = Contribution margin/Sales

= 135,000/270,000

= 50%

Break even point = Fixed Expenses/Contribution margin per unit

= 150,000/10

= 15,000 Units

Break Even point in Sales = Fixed Expenses/CM ratio

= 150,000/50%

= $300,000

2.Calculation of increase/(Decrease) in income;

Increase in Contribution margin = 86,000*50% = $43,000

Less: Additional Advertising $6,300

Increase in operating Income = $36,700

3.calculation of Income/Loss

Sales 27,000*18 = $486,000

Less: variable Expenses 27,000*10 = $270,000

Contribution Margin= $216,000

Less: Fixed Expenses = $150,000

Less: additional Advertising = $31,000

Operating income/(loss) = $35,000

4.Target profit = $4,300

Add: Fixed Costs = $150,000

Required Contribution Margin = $154,300

Contribution Margin per Unit = 10-0.4 = $9.60

Number of Units required to be sold = 154,300/9.60

=16,073 units

5.a.New CM ratio = (20-7)/20

= 65%

New Break Even point = (150,000+56,000)/65%

= $316,923

In Units = 206,000/13

= 15,846.15 units

b.Contribution format income Statement:

Automated Operations

Not Automated

Sales 20,300*20

406,000

406,000

Less: Variable Expenses

142,100

203,000

Contribution Margin

263,900

203,000

Less; Fixed Expenses

206,000

150,000

Net Operating Income

57,900

53,000

Yes, the company should automate its operations if it expects to sell 20,300 units